Circular No. 9107

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://www.cfa.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

COMET C/2010 A4 (SIDING SPRING) On Jan. 12, G. J. Garradd reported his discovery of an apparently

asteroidal object on CCD images taken with the 0.5-m Uppsala Schmidt telescope (discovery observation tabulated below); after posting on the Minor Planet Center’s ‘NEOCP’ webpage, Garradd later reported the object being elongated ∼ 10′′ in p.a. 150^{o} but with no clear tail in four 60-s stacked exposures. Other CCD astrometrists have also commented on the object’s cometary appearance, including W. H. Ryan (Magdalena Ridge Observatory, 2.4-m f/8.9 reflector, Jan. 13.35–13.38 UT), who noted a tail in p.a. ∼ 325^{o} in R-band images. E. Guido, Castellammare di Stabia, Italy, reports that twenty co-added 120-s unfiltered exposures taken by G. Sostero and himself remotely with a 0.25-m f/3.4 reflector located near Mayhill, NM, U.S.A., on Jan. 13.4 show a diffuse coma nearly 8′′ in diameter with a very weak central condensation. S. Foglia writes that thirty stacked 60-s images taken by R. Holmes (Ashmore, IL, U.S.A., 0.61-m reflector, Jan. 13.5) show the object to be diffuse.

2010 UT  α_{2000}  δ_{2000}  Mag. Observer
Jan. 12.72144 11^{h}02^{m}04.^{s}71 −27^{o}10′15′′.5 18.7 Garradd

Further astrometry, the following preliminary parabolic orbital elements, and an ephemeris are given on MPEC 2010-A76.

T = 2010 Oct. 8.813 TT  ω = 272.^{o}446
Ω = 346.731 }2000.0
q = 2.70743 AU  i = 96.846

COMET P/2010 A1 (HILL) Additional astrometry has shows this comet (cf. IAUC 9104) to be of

short period; elements from MPEC 2010-A50:

T = 2009 Aug. 9.353 TT  ω = 13.^{o}869
e = 0.55136  Ω = 47.080 }2000.0
q = 1.94218 AU  i = 10.265
a = 4.32903 AU n^{o} = 0.109426 P = 9.01 years
2010 January 14  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2010 CBAT  Daniel W. E. Green